OpenStudy (shiree19):

A true-false test has 12 questions. What is the probability of guessing the correct answers to all of the questions? Choices: A - 1/4096 B - 1/144 C - 1/24 D - 1/14 I choose C because 12 * 2 = 24 = 1/24 The correct answer is A - 1/4096 How do you get that?

OpenStudy (anonymous):

(1/2)^12 = 1/4096 like flipping a coin 12 times and getting heads each time

OpenStudy (shiree19):

I still don't understand. What does ^12 mean? Do I do 1 / 2 = 0.5

OpenStudy (anonymous):

a^12 = a x a x a...twelve times, a to the 12th power

OpenStudy (shiree19):

ok. a = 2. I get it now. Thanks so much! 2^12 = 4096 = 1/4096

OpenStudy (anonymous):

Very good!